You buy a bond for $1000 today that promises interest of $50 in one year plus the return of your principal. However, the probability that the company will default and not pay you either interest nor repay your principal is 1 percent. The expected return on the bond is ____ percent.
1
Expert's answer
2021-08-11T09:18:36-0400

Solution:

The expected value of a bond after one year = Sum (Return x Probability)

Where: Return = Expected return

            Probability = Expected probability

Probability of 0.01 will yield $0 expected returns

Probability of 0.99 will yield $1,050 expected returns

The expected value of a bond after one year = (0 "\times" 0.01) + (1050"\times" 0.99) = 0 + 1039.5 = "\$1,039.50"

The expected value of a bond after one year = "\$1,039.50"

 

Therefore, the expected return of a bond ="\frac{1039.50 - 1000}{1000} = 0.0395\times 100\% = 3.95 \%"